TMS Meeting Minutes, January 20, 2022

Attendees: Andrew, Breann, Christian, Heejoo, John, Olga, Matthieu, Trenton, Yang (Zoom Meeting)

Updates on Code, Testing Shutter:
  • Yang fixed the extraneous logging code to limit data logging to only necessary files
  • Heejoo also removed junk folders from the mountain machine
  • Heejoo ran a 21 hour data acquisition loop with no issues encountered with the shutter control
  • Etalon team now planning to move on to software testing to find the source of the errors
  • Though all fibers were unplugged earlier today, the team on the mountain was successful in reconnecting all fibers in preparation for tonight's observation
  • Please don't run any software tests until further notification
  • Yang still working on getting the missing channel code back into the updated TMS code (may be released tomorrow or early next week)
  • Heejoo to find time for a daytime test after code is released
  • For tonight's test, planning to run the old code with the correct channels
Olga's readme regarding procedure for last non/active-TMS test:
  • In the last test, TMS was active on Dx and not running on Sx
  • Christian identified a field that meets the criteria of our test for tonight
  • Olga wondering if we want to run the same sort of test, or a different test procedure
  • Olga has a question about the reference to be used for tonights testing
  • It is possible that this reference will be from the last TMS run on Dec. 13th, 2021
  • This will be the first night open after a couple nights of being closed
  • John notes that we should run as telescope@robs
  • Jenny will be observing, but Olga will be around
  • Andrew wondering if it's possible to do a quick continuity check to ensure all fibers are plugged in properly
  • The night starts with a quick AO observation to checkout, as soon as it's done moving to LBC's
Mountain Visit:
  • Mountain visit planned for next few weeks
  • First task, ensure that the collimator mounts are solid and not rocking
  • Then, installing fiber switch to enable multi-mode operation and more channels
  • Heejoo has not been able to find the exact same MPO casette that we have right now
  • Andrew suggests it would be best to contact Mark at Etalon about this before we go ahead
  • It is likely that the system is similar to standard telescom equipment, except for the fact that the fiber tips are angled and specialized for this system
  • Then, taking nice, high quality images for papers, publications, reports, etc.
  • Next, planning to count how many spare fibers we have and what kind of tip do these fibers have
  • If we find a very nice fiber with the correct tip but the fiber is broken, we can consider splicing fibers
Thermal Modeling:
  • Trenton is building a thermal model of the LBT primary mirror using Ansys software
Sensitizer:
  • Andrew talking to Ron H. about his Sensitizer code
  • Software able to extract FWHM and ellipiticity from a grid of differences on the field
  • Able to construct a least squares best fit to obtain gradients across the field
  • Sensitizer is an engine that plugs in to Zemax using a fifth-order model
  • Christian is focused mostly on getting the optical model updated with Sensitizer inputs

-- TrentonBrendel - 20 Jan 2022
Topic revision: r1 - 20 Jan 2022, TrentonBrendel
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback